&nb

                                                United States                                                            Canada

              Lumberjack shirts         Lumberjack boots                    Lumberjack shirts        Lumberjack boots

       A                   12                                     0                                                    6                                     0

       B                     3                                     3                                                    3                                     3

       C                      0                                    4                                                    0                                     6

Which country has absolute advantage in the production of shirts, in the production of boots? Why?

    a. Canada; United States                                         b. United States; Canada

12. In the United States, what is the opportunity cost of one shirt, and of one pair of boots? Show your calculation of domestic costs .Show costs calculation and circle the correct answer

a. 4 boots; 1/3 of shirt         b. 1/3 boots; 3 shirts                c.1 boot; 3 shirts            d. ¼ of boots; 1 shirt      

In Canada, what is the opportunity cost of one shirt, and of one pair of boots? Show your calculation

a. 1 boot; 1 shirt                    b. 3 boots; 3 shirts                   c. 1/3of boots; 1/3 of shirt
Answer #1


Question 11

Using same amount of resources in the production of lumberjack shirts, United States can produce 12 lumberjack shirts while Canada can produce 6 lumberjack shirts.

As United States can produce greater quantity of shirts relative to Canada using same amount of resources, US has absolute advantage in the production of lumberjack shirts.

Similarly,

Using same amount of resources in the production of lumberjack boots, United States can produce 4 lumberjack boots while Canada can produce 6 lumberjack boots.

As Canada can produce greater quantity of boots relative to United States using same amount of resources, Canada has absolute advantage in the production of lumberjack boots.

Hence, the correct answer is the option (b).

Question 12

(a)

Using all its resources, United States can produce a maximum of 12 shirts or 4 boots.

So,

Opportunity cost of one shirt = maximum number of boots that can be produced/maximum number of shirts that can be produced

Opportunity cost of one shirt = 4/12 = 1/3

The opportunity cost one shirt is 1/3 boots

Opportunity cost of one boot = maximum number of shirts that can be produced/maximum number of boots that can be produced

Opportunity cost of one boot = 12/4 = 3

The opportunity cost one boot is 3 shirts.

Hence, the correct answer is the option (b).

(b)

Using all its resources, Canada can produce a maximum of 6 shirts or 6 boots.

So,

Opportunity cost of one shirt = maximum number of boots that can be produced/maximum number of shirts that can be produced

Opportunity cost of one shirt = 6/6 = 1

The opportunity cost one shirt is 1 boots

Opportunity cost of one boot = maximum number of shirts that can be produced/maximum number of boots that can be produced

Opportunity cost of one boot = 6/6 = 1

The opportunity cost one boot is 1 shirt.

Hence, the correct answer is the option (a).